OpenStudy (anonymous):

please help Provide an example of a linear equation and explain how you could go about finding a few solutions for this particular linear equation. If you can, take this a step further by explaining how you could then use these solutions to graph the linear equation on the coordinate plane.

OpenStudy (anonymous):

2x+3y=6 is a linear equation to find points to plot the easiest way would be to set each x and y to 0 2(0)+3y=6 y=2 2x+3(0)=6 x=3

OpenStudy (anonymous):